OpenStudy (anonymous):

Square root -80/ square root -10

OpenStudy (agreene):

\[\sqrt{-80}/\sqrt{-10}\] \[=(4\sqrt{5}*i)/(\sqrt{10} *i)\]\[=2\sqrt{2}\] I skipped some simplification steps, but they are just from the laws of radicals.

OpenStudy (anonymous):

where did the square root 2 come from

OpenStudy (agreene):

sqrt(5)/sqrt(10) = 1/2 sqrt(2)

OpenStudy (anonymous):

so u flipped it

OpenStudy (agreene):

No, I just simplified the radicals to complex format, and then simplified from there, the 1/2 goes away because of the 4 in the numerator (4sqrt(5)) and 4* 1/2 = 2 and the i's go away because i/i = 1

OpenStudy (anonymous):

this makes no sense :(